OpenStudy (anonymous):

Divide 35b5 + 20ab3 + 20a2b2 by 5b2.

OpenStudy (anonymous):

Determine the independent and dependent variables, and then identify what the slope of a graph of this data represents. For every increase of 2.5 mph in the motorboat's speed, the parasail can fly 1 foot higher. For every increase of 0.4 mph in the motorboat's speed, the parasail can fly 5 feet higher. For every increase of 1 mph in the motorboat's speed, the parasail can fly 2.5 feet higher. For every increase of 2 mph in the motorboat's speed, the parasail can fly 2.5 feet higher.

OpenStudy (unklerhaukus):

\[\frac{35b^5 + 20ab^3 + 20a^2b^2}{5b^2}\\ =\frac{35b^5}{5b^2} + \frac{20ab^3}{5b^2} + \frac{20a^2b^2}{5b^2}\\=\dots\]

OpenStudy (unklerhaukus):

cancel common factors for each term, \[\frac{35b^5}{5b^2}= \frac{(\cancel{5}\cdot7)b^{\cancel 53}}{\cancel{5b^2}}=7b^3\]

OpenStudy (unklerhaukus):

what do you get when you cancel the factors of the second term @devonallison12 ? 20ab^3 / 5b^2 = ?

OpenStudy (unklerhaukus):

do you understand how i have broken up the big fraction into three terms , and canceled factors in the first term (that are common to the numerator and the denominator) @devonallison12 ?

OpenStudy (anonymous):

whats the answer x2- 8x + 16